### v2

# On Supportedness in Multi-Objective Combinatorial Optimization

Abstract
This paper addresses an inconsistency in various definitions of supported non-dominated points within multi-objective combinatorial problems (MOCO). MOCO problems are known to contain supported and unsupported non-dominated points, with the latter typically outnumbering the former. Supported points are, in general, easier to determine, can serve as representations, and are used in two-phase methods to generate the entire non-dominated point set. Despite their importance, several different characterizations for supported efficient solutions (and supported non-dominated points) are used in the literature. While these definitions are equivalent for multi-objective linear problems, they can yield different sets of supported non-dominated points for MOCO problems. We show by an example that these definitions are not equivalent for MOCO or general multi-objective optimization problems. Moreover, we analyze the structural and computational properties of the resulting sets of supported non-dominated points. These considerations motivate us to summarize equivalent definitions and characterizations for supported efficient solutions and to introduce a distinction between supported and weakly supported efficient solutions.
Keywords: multi-objective integer linear programming, multi-objective combinatorial optimization, supported efficient solutions, weakly supported efficient, weighted sum scalarization
1 Introduction
Multi-objective optimization (MOO) is concerned with optimizing multiple conflicting objectives simultaneously. For a comprehensive introduction to multi-objective optimization, we refer to Steuer, (1986) and Ehrgott, (2005).
Definition 1.1
A multi-objective optimization problem (MOO) is defined as
$\min_{x\in\mathcal{X}}f(x)=\left(f_{1}(x),\ldots,f_{p}(x)\right)^{\top},$ (MOO)
where $f\colon\mathds{R}^{n}\rightarrow\mathds{R}^{p}$ is a vector-valued objective function composed of $p\geq 2$ real-valued objective functions $f_{k}\colon\mathds{R}^{n}\rightarrow\mathds{R}$ for $k\in\{1,\ldots,p\}$ and $\mathcal{X}\subseteq\mathds{R}^{n}$ denotes the set of feasible solutions. We call $\mathds{R}^{n}$ the decision space and $\mathds{R}^{p}$ the objective space. The image of the feasible set
$\mathcal{Y}\coloneqq f(\mathcal{X})=\{f(x)\colon x\in\mathcal{X}\}\subseteq\mathds{R}^{p}$
is called the set of feasible outcome points in objective space.
The non-negative orthant of $\mathds{R}^{p}$ is denoted by $\mathbb{R}_{\geqq}^{p}:=\left\{x\in\mathbb{R}^{p}:x\geqq 0\right\}$ and analogously its interior $\mathds{R}_{>}^{p}$ and $\mathds{R}_{\geq}^{p}:=\left\{x\in\mathds{R}^{p}:x\geqq 0,x\neq 0\right\}$. In our notation, we also use the Minkowski sum and the Minkowski product of two sets $A,B\subseteq\mathds{R}^{p}$, which are defined as $A+B\coloneqq\{a+b\colon a\in A,b\in B\}$ and $A\cdot B\coloneqq\{a\cdot b\colon$ $a\in A,b\in B\}$, respectively. For a closed set $C$, we define the boundary of $C$ by $\partial C$.
The objective functions are typically assumed to be conflicting, implying the absence of an ideal solution that minimizes all objectives simultaneously. Instead, we rely on the Pareto concept of optimality, based on the component-wise order in $\mathds{R}^{p}$.
Definition 1.2
Let $y^{1},y^{2}\in\mathbb{R}^{p}$. We write:
  • $y^{1}\leqq y^{2}$ if $y_{i}^{1}\leq y^{2}_{i}$ for $i=1,\ldots,p$,
  • $y^{1}\leq y^{2}$ if $y^{1}\leqq y^{2}$ but $y^{1}\neq y^{2}$ and
  • $y^{1}<y^{2}$ if $y_{i}^{1}<y^{2}_{i}$ for $i=1,\ldots,p$.
Definition 1.3
A feasible solution $x^{*}\in\mathcal{X}$ is called efficient if there is no $x\in\mathcal{X}$ such that $f(x)\leq f\left(x^{*}\right)$. The image $f(x^{*})$ is then called a non-dominated point. The set of efficient solutions is denoted by $\mathcal{X}_{E}\subseteq\mathcal{X}$ and the set of non-dominated points by $\mathcal{Y}_{N}\subseteq\mathcal{Y}$.
A feasible solution $\hat{x}\in X$ is called weakly efficient if no $x\in\mathcal{X}$ exists with f(x)<f(x^)f(x)<f($\hat${x}). The image f(x^)f($\hat${x}) is then called a weakly non-dominated point. Furthermore, a feasible solution $x^{*}\in\mathcal{X}$ is called properly efficient (in Geoffrion’s sense (Geoffrion,, 1968)), if it is efficient and if there is a real number $M>0$ such that for all $k\in\{1,\ldots,p\}$ and $x\in\mathcal{X}$ satisfying $f_{k}(x)<f_{k}({x}^{*})$ there exists an index $j$ such that $f_{j}({x}^{*})<f_{j}(x)$ such that
$\frac{f_{k}({x}^{*})-f_{k}(x)}{f_{j}(x)-f_{j}({x}^{*})}\leq M.$
The image ${y}^{*}=f({x}^{*})$ is called properly non-dominated.
In the following, we assume that the multi-objective optimization problem has at least one non-dominated point, i.e., $\mathcal{Y}_{N}\neq\varnothing$. The polyhedron $\mathcal{Y}^{\geqq}:=\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$ is called the upper image of $\mathcal{Y}$.
Scalarization approaches are the most common solution techniques in multi-objective optimization, which rely on replacing the multi-objective problem by a sequence of single-criteria optimization problems. In the weighted-sum scalarization a convex combination (with weights $\lambda_{i}$) of the objective functions is optimized over the feasible set $\mathcal{X}$. Then, the set of normalized weighting vectors (or weight space) is defined as the set $\Lambda_{p}=\{\lambda\in\mathds{R}^{p}_{>}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ or $\Lambda_{p}^{0}=\{\lambda\in\mathds{R}^{p}_{\geq}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ if weights equal to zero are included.
Definition 1.4
The weighted-sum scalarization of (MOO) with weighting vector $\lambda\in\Lambda_{p}$ or $\lambda\in\Lambda_{p}^{0}$ is defined as the parametric program
$P_{\lambda}\coloneqq\min_{x\in\mathcal{X}}\;\lambda^{\top}f(x)$ ($P_{\lambda}$)
Theorem 1.5 (Ehrgott, 2005)
If $\lambda\in\Lambda_{p}^{0}$, every optimal solution of $P_{\lambda}$ is a weakly efficient solution of (MOO). Moreover, every optimal solution of $P_{\lambda}$ problem is properly efficient for (MOO), if $\lambda\in\Lambda_{p}$.
A specific, well-studied class of MOO are multi-objective linear programs (MOLP) and multi-objective integer linear programs (MOILP).
Definition 1.6
A multi-objective linear optimization problem (MOLP) is defined by
$\min_{x\in\mathcal{X}}y(x)=\left(y_{1}(x),\ldots,y_{p}(x)\right)^{\top}=Cx,$ (MOLP)
with $\mathcal{X}\coloneqq\{x\in\mathds{R}^{n}\colon Ax\leqq b,x\geqq 0\}$. Thereby, the rows $c^{k}$ ($k=1,\ldots,p$) of the cost matrix $C\in\mathbb{R}^{p\times n}$ contain the coefficients of the $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and the matrix $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m}$ formulates the $m$ linear constraints.
Definition 1.7
If all decision variables are further restricted to the set of integers, i.e., $x\in\mathds{Z}$, we obtain a multi-objective integer linear problem (MOILP).
$\min_{x\in\mathcal{X}}y(x)=\left(y_{1}(x),\ldots,y_{p}(x)\right)^{\top}=Cx,$ (MOILP)
with $\mathcal{X}\coloneqq\{x\in\mathds{Z}^{n}\colon Ax\leqq b,x\geqq 0\}$ and all other elements as described in (MOLP).
Theorem 1.8 (Isermann, 1974)
Considering MOLP problems, for every efficient solution $\bar{x}\in\mathcal{X}_{E}$ there exists a $\lambda\in\Lambda_{p}$ such that $\bar{x}$ is a optimal solution of the weighted sum scalarization ($P_{\lambda}$).
Hence, for a MOLP, the efficient set $\mathcal{X}_{E}$ and the union over all sets of optimal solutions of $P_{\lambda}$ for $\lambda\in\Lambda_{p}$ are identical. Thus, in multi-objective linear programs (MOLP) every non-dominated point lies on the non-dominated frontier, which can be defined in the following way.
Definition 1.9 (Hamacher et al., (2007))
The non-dominated frontier is the set
$\Bigl{\{}y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\Bigr{\}}.$
Note that Hamacher uses the wording efficient frontier instead of the non-dominated frontier. The non-dominated frontier equals the non-dominated set of $\operatorname{conv}(\mathcal{Y}_{N})$ and can be characterized as the set of the maximal non-dominated faces of the upper image $\mathcal{Y}^{\geqq}$ (Ehrgott,, 2005). Hereby, a non-dominated face $F\subseteq\mathcal{Y}^{\geqq}$ is a face of the upper image $\mathcal{Y}^{\geqq}$ such that every point on $F$ is non-dominated with respect to $\mathcal{Y}^{\geqq}$. Similarly, $F\subseteq\mathcal{Y}^{\geqq}$ is called a weakly non-dominated face if all its points are weakly non-dominated. A face $F$ is called maximally non-dominated if there is no other non-dominated face $G$ of $\mathcal{Y}^{\geqq}$ such that $F$ is a proper subset of $G$. This would imply that the dimension of $G$ is greater than the dimension of $F$. The preimage $F_{\mathcal{X}}$ of a maximally non-dominated face $F_{\mathcal{Y}}$ of $\mathcal{Y}^{\geqq}$, i.e., all solutions whose image lies in $F_{\mathcal{Y}}$, is denoted as the maximally efficient face. Note that $F_{\mathcal{X}}$ might not be a face of the feasible set, the polyhedron $\mathcal{X}$, since multiple feasible solutions (located on different faces of $\mathcal{X}$) could map to the same non-dominated point $y\in\mathcal{Y}$.
$y_{3}$$y_{1}$$y_{2}$
Figure 1: Illustration of dominance relations in faces: the red triangular face represents a maximal non-dominated face; its red boundary edges are non-dominated faces; and the light blue faces are weakly non-dominated.
Figure 1 illustrates the geometric structure of non-dominated and weakly non-dominated faces for a continuous multi-objective minimization problem, where the feasible region consists of a cube with a cut at its lower-left front corner. The red triangular region corresponds to a maximal non-dominated face, meaning no point in this region is dominated by another feasible point, and this face cannot be extended to a larger non-dominated face within the feasible set. The bold red edges of this triangle represent lower-dimensional non-dominated faces, while the light-blue shaded cube faces indicate weakly non-dominated faces, which are optimal in some but not all directions.
In contrast to multi-objective linear programs (MOLP), where every non-dominated point lies on the non-dominated frontier, and thus on the convex hull of the upper image, MOILPs or MOCO problems may contain non-dominated points that are located in the interior of the upper image, i.e., in $\operatorname{int}(\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq}))$ and cannot be generated through any weighted-sum scalarization.
Definition 1.10
Formally a multi-objective combinatorial optimization (MOCO) problem is defined as
$\min_{x\in\mathcal{X}}y(x)=\left(y_{1}(x),\ldots,y_{p}(x)\right)^{\top}=C\,x,$ (MOCO)
where $\mathcal{X}\coloneqq\{x\in\{0,1\}^{n}\colon Ax=b\}$, with the cost matrix $C\in\mathbb{Z}^{p\times n}$ containing the rows $c^{k}$ of coefficients of $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and $A\in\mathbb{Z}^{m\times n},b\in\mathbb{Z}^{m}$ describing the $m$ constraints. The constraints define combinatorial structures such as paths, trees, or cycles in a network or partitions of a set.
This definition includes various problems such as multi-objective spanning trees, shortest paths, knapsack and assignment problems.
Two major challenges are associated with MOCO problems. Firstly, they are theoretically more challenging than their single-objective counterparts, falling into the class of computationally intractable problems (Ehrgott,, 2005). Secondly, in MOCO problems, non-dominated points located in the interior of the upper image can occur, which typically outnumber the set of solutions on the non-dominated frontier (see, e.g., Visẽe et al., 1998). Hence, a distinction between different classes of efficient solutions is required. An intuitive definition of supportedness is the following.
Definition 1.11
  1. 1.
    An efficient solution is called supported efficient solution if it is an optimal solution to the weighted sum scalarization $P_{\lambda}$ for $\lambda\in\Lambda_{p}$, i.e., an optimal solution to a single objective weighted-sum problem where the weights are strictly positive. Its image is called supported non-dominated point; we use the notation $\mathcal{X}_{S}$ and $\mathcal{Y}_{S}$ for the supported efficient solution set and supported non-dominated point set, respectively. For discrete problems, supported non-dominated points are located on the non-dominated frontier, i.e., located on the union of the maximal non-dominated faces of the upper image.
  2. 2.
    Extreme supported efficient solutions are those solutions whose image lies on the vertex set of the upper image. Their image is called an extreme supported non-dominated point.
  3. 3.
    Unsupported efficient solutions are efficient solutions that are not optimal solutions of $P_{\lambda}$ for any $\lambda\in\Lambda_{p}^{0}$. Unsupported non-dominated points lie in the interior of the upper image.
Figure˜2 illustrates supported extreme, supported, and unsupported non-dominated points, as well as the upper image in the bi-objective case.
$y_{1}$$y_{2}$$\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{2}_{\geqq})$extreme supported non-dominatedsupported non-dominatedunsupported non-dominateddominated
Figure 2: Illustration of the upper image $\mathcal{Y}^{\geqq}=\operatorname{conv}(\mathcal{Y})+\mathds{R}^{2}_{\geqq}$ and the different solution types.
As unsupported non-dominated points cannot be computed through a weighted sum scalarization, different scalarization techniques are required that are often computationally more expensive.
The determination of supported non-dominated points has gained attention for several reasons. First, these points are generally easier to determine than the unsupported non-dominated points and, moreover, the unsupported non-dominated points in MOCO typically outnumber the supported ones, as observed in Studies as Visẽe et al., (1998).
Second, they can serve as a foundation for the second phase of two-phase methods, which aims to generate the entire non-dominated point set using information derived from the supported non-dominated points (Pasternak and Passy, 1972; Visẽe et al., 1998; Hamacher et al., 2007; Przybylski et al., 2008; Przybylski et al., 2010b ; Eusébio and Figueira, 2009a ; Dai and Charkhgard, 2018,…). Recently, a computational study Sayin, (2024) showed that the set of supported non-dominated points can be used as a high-quality representation in multi-objective discrete optimization problems, focusing specifically on binary knapsack and assignment problems.
Several studies focus on identifying or analyzing the supported or extreme supported point set across various combinatorial problems, including bi- and multi-objective integer network flow problems (Eusébio and Figueira, 2009c, ; Medrano and Church,, 2015; Raith and Sedeño-Noda,, 2017; Könen and Stiglmayr,, 2025, 2023), bi- and multi-objective minimum spanning trees (Sourd et al.,, 2006; da Silva and Clí­maco,, 2007; Correia et al.,, 2021), bi- and multi-objective shortest path problems (Edwin Romeijn and Smith,, 1999; Sedeño-Noda and Raith,, 2015; Medrano and Church,, 2014), bi- and multi-objective combinatorial unconstrained problems (Bökler,, 2018; Schulze et al.,, 2019), bi- and multi-objective knapsack problems (Visẽe et al.,, 1998; Argyris et al.,, 2011; Schulze,, 2017), and bi- and multi-objective assignment problems (Tuyttens et al.,, 2000; Gandibleux et al.,, 2003; Przybylski et al., 2010b, ). Additional work on general MOCO problems can be found in Gandibleux et al., (2001); Jesus, (2015); Sayin, (2024), among others.
Beyond MOCO problems, there is research on the identification of supported solutions in multi-objective mixed integer problems (Özpeynirci and Köksalan,, 2010; Pettersson and Ozlen,, 2019; Bökler et al.,, 2024) or addressing supportedness in non-convex problems (The Luc,, 1995; Liefooghe et al.,, 2014, 2015). Summarizing, supported points are often easier to determine, can serve as high-quality representations, and can be used in two-phase methods to generate the entire non-dominated point set.
However, despite their significance, several alternative definitions (as in Definition˜1.11) and characterizations for supportedness are used. These different definitions differ in the literature and sometimes even within a single publication. The following chapter proves that these definitions are not equivalent in the case of discrete or more general optimization problems. As a result, they generate different sets of supported efficient solutions and, consequently, different supported non-dominated point sets with varying properties.
This motivates us to distinguish between supported efficient solutions and weakly supported efficient solutions. An efficient solution is denoted as weakly supported if it is an optimal solution of a weighted sum scalarization with non-negative weights. In contrast to the definition of supportedness, weakly supportedness allows single weights to have a value of zero. Note that the terminology weakly supported efficient solutions, despite the fact that these solutions are still efficient, is motivated by geometric considerations: any weakly supported non-dominated point that is not a supported non-dominated point lies on a weakly non-dominated face, but not on a non-dominated face. In contrast, supported points necessarily lie on non-dominated faces.
This work focuses on the theoretical understanding of supportedness in MOCO problems. In particular, we analyze the structural properties of supported and weakly supported non-dominated points and highlight important distinctions through counterexamples. However, further investigation into the supportedness of general MOO problems remains sparse and presents an important area for future research. A recent study in (Chlumsky-Harttmann,, 2025) provided a first approach toward a categorization of supportedness definitions for general MOO problems.
The remainder of the paper is structured as follows. In Section˜2 the different definitions and characterizations of supportedness found in the literature are presented. A counterexample is given that shows that these definitions are not equivalent for MOILPs and MOCO problems. Additionally, it discusses the distinction between supported and weakly supported non-dominated points. Finally, Section˜3 concludes the findings and outlines directions for future research.
2 Supported and Weakly Supported Non-dominated Points in MOCO problems
Several definitions and characterizations despite Definition˜2.1 of supported non-dominated points exist in the literature and sometimes are even inconsistent within a single publication.
Definition 2.1 (Conflicting Definitions of Supportedness)
A point $y^{\prime}\in\mathcal{Y}$ is called supported
  • (1)
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ lies on the boundary of the upper image $\mathcal{Y}^{\geqq}\coloneqq\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$, i.e., $y^{\prime}\in\mathcal{Y}_{N}\cap\partial\mathcal{Y}^{\geqq}$ (Eusébio and Figueira, 2009b, ; Eusébio and Figueira, 2009c, ; Liefooghe et al.,, 2014, 2015; Correia et al.,, 2021). The corresponding set of supported non-dominated points is denoted by $\mathcal{Y}_{S\partial}$.
  • (2)
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ is located on the non-dominated frontier defined as the set $\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$ (Hamacher et al.,, 2007), the set supported non-dominated points according to this definition is denoted by $\mathcal{Y}_{SNF}$.
  • (3)
    if $y^{\prime}$ is the image of a supported efficient solution, which are those efficient solutions that can be obtained as optimal solutions of a weighted sum scalarization with weights strictly greater zero, i.e. $y$ is an image of a solution $x\in\arg\min P_{\lambda}$ for a $\lambda\in\Lambda_{p}$ (Visẽe et al.,, 1998; Ehrgott,, 2005; Raith and Ehrgott,, 2009; Przybylski et al., 2010a, ; Argyris et al.,, 2011; Raith and Sedeño-Noda,, 2017). The set of supported non-dominated points obtained with respect to this definition is denoted by $\mathcal{Y}_{S\lambda}.$
Note, that some publications adopt the definition related to $\mathcal{Y}_{S\lambda}$ but allowing $\lambda\in\Lambda_{p}^{0}$, provided that $y$ is non-dominated, see, e.g. Gandibleux et al., (2001); Liefooghe et al., (2015). Regarding definition $\mathcal{Y}_{SNF}$Hamacher et al., (2007) uses the wording efficient frontier instead of the non-dominated frontier. There also exist definitions based on convex combinations of the non-dominated points as given in Özpeynirci and Köksalan, (2010).
Isermann, (1974) showed that for a MOLP, the set of non-dominated efficient solutions coincides with the set of optimal solutions of the weighted sum method with weights strictly greater than zero, i.e., in  MOLP, the weighted sum scalarization achieves completeness when varying $\lambda$ within $\Lambda_{p}$. Consequently, for a  MOLP, it holds $\mathcal{Y}_{S\partial}=\mathcal{Y}_{SNF}=\mathcal{Y}_{S\lambda}$.
Theorem 2.2
For a MOLP it holds $\mathcal{Y}_{S\partial}=\mathcal{Y}_{SNF}=\mathcal{Y}_{S\lambda}$.
Proof.
According to ˜1.8, any non-dominated point $y^{\prime}\in\mathcal{Y}$ in MOLP can be obtained as the image of an optimal solution $x\in\arg\min P_{\lambda}$ with $\lambda\in\Lambda_{p}$, and with ˜1.5 we obtain $\mathcal{Y}_{S\lambda}=\mathcal{Y}_{N}$. Any point $\hat{y}\in\mathcal{Y}_{S\partial}$ or $\hat{y}\in\mathcal{Y}_{SNF}$ is non-dominated per definition and therefore $\hat{y}\in\mathcal{Y}_{S\lambda}.$ It follows
$\mathcal{Y}_{S\lambda}\supseteq\mathcal{Y}_{S\partial}\text{ and }\mathcal{Y}_{S\lambda}\supseteq\mathcal{Y}_{SNF}.$
[…]
$\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{S\partial}\text{ and }\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{SNF}.$
This concludes $\mathcal{Y}_{S\partial}=\mathcal{Y}_{SNF}=\mathcal{Y}_{S\lambda}$. ∎
This equivalence does not hold for the discrete or the non-linear case. In the literature, supported non-dominated points for MOCO or MOILP are often characterized as non-dominated points that lie on the boundary of the upper image and that they only lie on the maximally non-dominated faces, i.e., the non-dominated frontier. In contrast, the unsupported are characterized as non-dominated points that lie in the interior of the upper image, e.g., (Eusébio and Figueira, 2009c, ; Przybylski et al., 2010a, ; Correia et al.,, 2021).
However, depending on the definition, unsupported non-dominated points may exist that lie on the boundary of the upper image or supported non-dominated points that lie on weakly non-dominated faces and thus do not lie on the non-dominated frontier. In this case, they cannot be obtained as optimal solutions of a weighted sum problem with weights strictly greater than zero. Suppose a supported non-dominated point is defined according to Definition˜2.1 (2) and (3), i.e., considering the sets $\mathcal{Y}_{S\lambda}$ and $\mathcal{Y}_{SNF}$. In that case, unsupported points may exist on the boundary of the upper image. Conversely, if a supported non-dominated point is defined according to (1), i.e., points contained in $\mathcal{Y}_{S\partial}$, there may exist supported non-dominated points on the boundary of the upper image which are not lying on the non-dominated frontier, i. e., which lie on weakly non-dominated faces. This inconsistency motivates us to develop new, consistent definitions of supported and weakly supported non-dominated points.
Definition 2.3 (Supported/Weakly Supported)
An efficient solution is called a weakly supported efficient solution if it is an optimal solution of a weighted-sum scalarization ${P}_{\lambda}$ for some weight $\lambda\in\Lambda_{p}^{0}$. Moreover, if the weight is strictly positive $\lambda\in\Lambda_{p}$, it is called supported efficient solution. The corresponding image is called weakly supported or supported (non-dominated) point, respectively.
Note that the terminology weakly supported efficient solutions, despite the fact that these solutions are still efficient, is motivated by geometric considerations: any weakly supported non-dominated point that is not a supported non-dominated point lies on a weakly non-dominated face, but not on a non-dominated face. In contrast, supported points necessarily lie on non-dominated faces.
Let $\mathcal{Y}_{S}$ and $\mathcal{Y}_{wS}$ be the set of all supported non-dominated points and the set of all weakly supported non-dominated points, respectively. Accordingly, $\mathcal{X}_{S}$ ($\mathcal{X}_{wS})$ is the set of all (weakly) supported efficient solutions. We denote the cardinality of $\mathcal{X}_{S}$ by $|\mathcal{X}_{S}|=\mathcal{S}$.
With these definitions, the following characterizations for MOCO problems hold: While weakly supported non-dominated points lie on the boundary of the upper image, supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces. The unsupported non-dominated points lie in the interior of the upper image.
Theorem 2.4
$\mathcal{Y}_{wS}=\mathcal{Y}_{S\partial}$
Proof.
We follow the proofs of Theorem 3.4 and Theorem 3.5 in Ehrgott, (2005) with slight modifications. To prove the equality of the two sets, we show that the subset relation holds in both directions.
  • $\mathcal{Y}_{wS}\subseteq\mathcal{Y}_{S\partial}$: Let $\hat{y}\in\mathcal{Y}_{wS}$, i.e., there exists a $\lambda\in\Lambda^{0}_{p}$ such that $\hat{y}$ is an image of an optimal solution of ($P_{\lambda}$) and $\hat{y}$ is non-dominated.
    Suppose that $\hat{y}\notin\mathcal{Y}_{S\partial}$, i.e., $\hat{y}\notin\partial\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Then $\hat{y}$ must lie in the interior of $\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Thus, there exists an $\varepsilon$-neighborhood $B(\hat{y},\varepsilon)$ of $\hat{y}$, defined as $B(\hat{y},\varepsilon):=\hat{y}+B(0,\varepsilon)\subset\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$, where $B(0,\varepsilon)$ is an open ball with radius $\varepsilon$ centered at the origin. Let $d\in\mathbb{R}^{p}_{>}$. Then we can choose some $\alpha\in\mathbb{R}$, $0<\alpha<\varepsilon$ such that $\alpha\,d\in B(0,\varepsilon)$. Now, $y^{\prime}=\hat{y}-\alpha\,d\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$ with $y_{k}^{\prime}<\hat{y}_{k}$ for $k=1,\dots,p$ and
    $\sum_{k=1}^{p}\lambda_{k}\,y_{k}^{\prime}<\sum_{k=1}^{p}\lambda_{k}\,\hat{y}_{k},$
    because at least one of the weights $\lambda_{k}$ must be positive. This contradiction implies the result.
  • $\mathcal{Y}_{S\partial}\subseteq\mathcal{Y}_{wS}:$ Let $\hat{y}\in\mathcal{Y}_{S\partial}\implies\hat{y}\in\partial\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$ and $\hat{y}$ is non-dominated, hence it exists a supporting hyperplane $\{y\in\mathds{R}^{p}:\lambda^{\top}y=\lambda^{\top}\hat{y}\}$ with $\lambda\in\mathds{R}^{p}\setminus\{0\}$ and it holds $\lambda^{\top}y\geq\lambda^{\top}\hat{y}$ for all $y\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Furthermore, by applying the separation theorem to the disjoint sets $\{y+d\colon d\in\mathds{R}^{p}_{>}\}$ and $\{\hat{y}-d^{\prime}\colon d^{\prime}\in\mathds{R}^{p}_{>}\}$, we obtain:
    $\lambda^{\top}(y+d-\hat{y})\geq 0\geq\lambda^{\top}(-d^{\prime})$
    for all $y\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$, and $d,d^{\prime}\in\mathds{R}^{p}_{>}.$ Choose $d^{\prime}=e_{k}+\varepsilon\,e$, where $e_{k}$ is the $k$-th unit vector and $e=(1,\ldots,1)^{\top}\in\mathds{R}^{p}$. With $\varepsilon>0$ arbitrary small we see that $\lambda_{k}\geq 0$ for all $k=1,\ldots,p$. Thus, $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ with $\lambda\in\Lambda^{p}_{0}$ and per definition non-dominated. It follows $\hat{y}\in\mathcal{Y}_{wS}$.
Note that although the following results are stated for MOCO, they can be directly transferred to MOILP.
Theorem 2.5
For MOCO problems, it holds $\mathcal{Y}_{S}=\mathcal{Y}_{SNF}.$
Proof.
[…]
$\mathcal{Y}_{S}\supseteq\mathcal{Y}_{SNF}:$ Any point $\hat{y}\in\mathcal{Y}_{SNF}$ is a minimal element of the set $\operatorname{conv}(\mathcal{Y}_{N})$ induced by the cone $\mathds{R}_{\geqq}^{p}$. Thus $\hat{y}$ is a non-dominated point for the set $\operatorname{conv}(\mathcal{Y}_{N})$. If we define the relaxation MOLP of the given MOCO problem by
$\min_{x\in\operatorname{conv}(\mathcal{X})}C\,x,$
it follows that $\hat{y}$, is an image of an optimal solution for the reduced MOLP. According to ˜1.8, there is a $\lambda\in\Lambda_{p}$ such that $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ and hence $\hat{y}\in\mathcal{Y}_{S}$. ∎
Consequently, Definition 2.1 (1) includes the weakly supported non-dominated points for general MOO problems, while Definitions (3) and (4) only contain the supported non-dominated points (in the discrete case). Hence, in the discrete case, the weakly supported non-dominated points lie on the boundary of the upper image while the supported ones lie only on the non-dominated frontier, i. e., on maximally non-dominated faces.
Example 2.6
To illustrate the geometrical properties of weakly supported non-dominated points, consider the outcome vectors in Figure˜3. The outcome vectors are partitioned into two layers based on their value in the third component. The non-dominated points $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$ all share a minimum value of $c_{3}=1$ in their third component (pink layer), while there are other non-dominated points having a value of 5 in the third component (blue layer).
Among them, only $y^{1},y^{2},$ and $y^{3}$ lie on the non-dominated frontier and can be obtained as optimal solutions of a weighted-sum scalarization with $\lambda\in\Lambda_{d}$. In contrast, $y^{4}$ does not lie on the non-dominated frontier but is still on the boundary of the upper image (on a weakly non-dominated face). It can be obtained as an optimal solution of a weighted-sum scalarization where some weights are zero, e.g., obtained as an optimal solution of $P_{\lambda}$ with $\lambda^{\top}=(0,0,1)^{\top}$, making $y^{4}$ a weakly supported but not a supported non-dominated vector.
$y_{3}$$y_{1}$$y_{2}$$y_{3}=5$$y_{3}=1$$y^{1}$$y^{2}$$y^{4}$$y^{3}$
$y_{1}$$y_{2}$$y_{3}=1$$y^{1}$$y^{2}$$y^{4}$$y^{3}$
Figure 3: Outcome space with non-dominated points $y^{1},\ldots,y^{4}$, each with the value of 1 in the third component and the two-dimensional projection of the plane $c_{3}=1$.
Theorem 2.7
Let $\mathcal{Y}_{S}$ and $\mathcal{Y}_{wS}$ be the sets of supported and weakly supported non-dominated points of a MOCO problem, respectively. Then, $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ and there exist instances where $\mathcal{Y}_{S}\subset\mathcal{Y}_{wS}$, i. e., the set of supported non-dominated points is a proper subset of the set of the weakly supported non-dominated points.
Proof.
The inclusion $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ holds per definition. Consider the outcome set $\mathcal{Y}=\{y^{1},y^{2},y^{3},y^{4}\}$ with $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$, which represents the pink layer of the outcome set in Example˜2.6, displayed in Figure 3. It is straightforward to construct an artificial MOILP problem with this outcome set.
All points $y^{i}$ with $i\in\{1,\ldots,4\}$ are weakly supported non-dominated points since their preimages are optimal solutions of a weighted sum problem $P_{\lambda}$ with $\lambda=(0,0,1)^{\top}$. Since $y^{1}$$y^{2}$ and $y^{3}$ are also supported non-dominated points, since their preimages are optimal solutions of the respective weighted sum problems $P_{\lambda^{1}}$ with $\lambda^{1}=(0.7,0.1,0.2)^{\top}$, $P_{\lambda^{2}}$ with $\lambda^{2}=(0.4,0.4,0.2)^{\top}$, and $P_{\lambda^{3}}$ with $\lambda^{3}=(0.1,0.8,0.1)^{\top}$, respectively. The corresponding (projected) weight space decomposition is illustrated in Figure˜4. Note that, for example, $(6,4.2,1)^{\top}\in\operatorname{conv}(\mathcal{Y}_{N})\cap(y^{4}-\mathds{R}^{p}_{\geqq})$ implies $y^{4}\notin\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$. By ˜2.5, it follows that $y_{4}\notin\mathcal{Y}_{S}$, although $y^{4}\in\mathcal{Y}_{wS}$.
Thus, the set of weakly supported non-dominated points is $\mathcal{Y}_{wS}=\{y^{1},y^{2},y^{3},y^{4}\}$, while the set of supported non-dominated points is $\mathcal{Y}_{S}=\{y^{1},y^{2},y^{3}\}$. Hence, in this example, the set of supported non-dominated points is a proper subset of the set of weakly supported non-dominated points $\mathcal{Y}_{S}\subset\mathcal{Y}_{wS}$. ∎
[…]
Figure 4: Projected weight space decomposition to the upper image of $\mathcal{Y}=\{(2,9,1)^{\top},(3,6,1)^{\top},(8,3,1)^{\top},(6,5,1)^{\top}\}$ with $\lambda_{3}\coloneqq 1-\lambda_{2}-\lambda_{1}.$ The set of weighting vectors associated with a point $y\in\mathcal{Y}$ is given by $\Lambda(y)\coloneqq\{\lambda\in\Lambda^{0}_{p}\colon\lambda^{\top}y\leq\lambda^{\top}y^{\prime}\text{ for all }y^{\prime}\in\mathcal{Y^{\geq}}\}.$ For a comprehensive overview of the weight space decomposition, we refer to Przybylski et al., 2010a .
Note that in the bi-objective case $\mathcal{Y}_{S}=\mathcal{Y}_{wS}$ as the following lemma shows.
Lemma 2.8
Every weakly supported non-dominated point of a biobjective integer optimization problem is supported non-dominated.
Proof.
Let $\bar{y}=f(\bar{x})\in\mathds{R}^{2}$ be a weakly supported but not supported non-dominated point of a bi-objective integer optimization problem and $\bar{x}$ the corresponding preimage. Then there is a weighting vector $\lambda\in\Lambda_{0}$ such that $\bar{x}$ is an optimal solution of the weighted sum problem $P_{\lambda}$. Since $\bar{y}$ is not supported non-dominated, one of the components of $\lambda$ must be zero, w.l.o.g. let $\lambda=(1,0)^{\top}$.
Since $y$ is non-dominated, there does not exist a feasible outcome vector $y=(y_{1},y_{2})^{\top}$, with $y_{1}=\bar{y}_{1}$ and $y_{2}<\bar{y}_{2}$. Thus, $\bar{x}$ is also an optimal solution of the weighted sum problem $P_{\lambda^{\prime}}$ with $\lambda^{\prime}=(1-\varepsilon,\varepsilon)^{\top}$ for $\varepsilon>0$ sufficiently small, which makes $\bar{y}$ supported non-dominated. ∎
The clear distinction between the sets of supported non-dominated and weakly supported non-dominated solutions (Definition˜2.3) is also necessary as the corresponding problems may differ in their output time complexity. For instance, in the case of the minimum cost flow problem, it can be shown that supported efficient solutions can be determined in output-polynomial time, while this is not the case for weakly supported solutions, unless $\mathbf{P}=\mathbf{NP}$ (Könen and Stiglmayr,, 2025).
Theorem 2.9
The determination of all weakly-supported solutions for a MOCO problem with $p+1$ objectives is as hard as the determination of all non-dominated points for a MOCO with $p$ objectives.
Proof.
Assume an algorithm exists to determine all weakly supported non-dominated points for a given MOCO problem with $p+1$ objectives. Let $M_{p}$ be a MOCO with $p$ objectives. Suppose we add an artificial objective $c^{p+1}=0$ to our MOCO problem and denote it by $M_{p+1}$. In that case, we obtain a weakly efficient facet for $M_{p+1}$, where all non-dominated points for $M_{p}$ are weakly supported non-dominated points for $M_{p+1}$. Therefore, even the unsupported non-dominated points for $M_{p}$ are weakly supported non-dominated points for $M_{p+1}$ since they are part of the boundary of the upper image for $M_{p+1}$. Consequently, any algorithm that can determine all weakly supported non-dominated points for $M_{p+1}$ can determine all non-dominated points for $M_{p}$. ∎
While the above characterization of supported points holds for MOCO problems,it does not extend to general MOO problems, and it can be shown that in the case of general MOO problems it may hold $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. This distinction arises because the weighted sum scalarization for $\lambda\in\Lambda_{p}$ method is only capable of identifying properly supported non-dominated points in the sense of Geoffrion (Geoffrion,, 1968), as discussed in Ehrgott, (2005). Note that for MOCO problems every efficient solution is also properly efficient (Ehrgott,, 2005). However, for general MOO problems, the set of properly supported non-dominated points may form a strict subset of the non-dominated points on the non-dominated frontier, highlighting the limitations of the weighted sum approach in capturing the entire set of supported solutions in general MOO problems (Ehrgott,, 2005). The following example, similar to the ones in Boyd and Vandenberghe, (2004) or Ehrgott, (2005), illustrates this.
Example 2.10
Consider the following MOO problem
$\min\quad$ $x_{1}+2$
$\min\quad$ $x_{2}+2$
s.t. $x_{1}^{2}+x_{2}^{2}\leq 1$
Consider the points $y^{1}=(1,2)$ and $y^{2}=(2,1)$, which can be obtained with $\lambda^{1}=(0,1)$ and $\lambda^{2}=(1,0)$, respectively. It holds $y^{1},y^{2}\in\mathcal{Y}_{SNF}$ but $y^{1}$ and $y^{2}$ are not optimal solutions for $P_{\lambda}$ for any $\lambda\in\Lambda$ (Boyd and Vandenberghe,, 2004; Ehrgott,, 2005). Thus, $y^{1},y^{2}\notin\mathcal{Y}_{S}$. However, any point on the left lower boundary without the points $(1,2)$ and $(2,1)$ are both contained in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S}$. Therefore, in this example, $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. The example is illustrated in Figure˜5
$f_{2}$$f_{1}$$y^{2}-\mathds{R}^{2}_{\geqq}$$y^{1}-\mathds{R}^{2}_{\geqq}$$\mathcal{Y}$$y^{1}=(\begin{smallmatrix}1\\ 2\end{smallmatrix})$$y^{2}=(\begin{smallmatrix}2\\ 1\end{smallmatrix})$
Figure 5: Illustration of Example˜2.10 highlighting the points $y^{1},y^{2}$ the points that contains the points that are both in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S\lambda}$ (bold curve), which together with $y^{1}$ and $y^{2}$ gives the non-dominated frontier.
Theorem 2.11
For general MOO problems it holds $\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{SNF}$ and there exist instances where $\mathcal{Y}_{S\lambda}\subset\mathcal{Y}_{SNF}$.
Proof.
$\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{SNF}$ follows directly from the proof of ˜2.5. Furthermore, as demonstrated in Example˜2.10, there exist instances where $\mathcal{Y}_{S\lambda}\subset\mathcal{Y}_{SNF}$. ∎
The distinction between supported and weakly supported is sufficient for MOCO and MOILP problems. However, in the context of general MOO, it may be worthwhile to introduce a finer classification, distinguishing between properly supported, supported, and weakly supported non-dominated points. While this work primarily focuses on the supportedness in MOCO problems, a comprehensive analysis of supportedness in general MOO problems remains an important area for future research. Notably, recent advancements in this direction have been discussed in Chlumsky-Harttmann, (2025).
3 Conclusion
Some previous papers use inconsistent characterizations of supported non-dominated points for MOCO problems. Through counterexamples and theoretical analysis, this paper proves that these definitions, while being equivalent in the context of MOLP, diverge in MOCO problems, yielding distinct sets of supported non-dominated points with differing structural and computational properties. This emphasizes the need for precise and consistent definitions. Hence, this paper proposes the definition of weakly supported non-dominated points and establishes a clear distinction between weakly supported and supported non-dominated points. This is particularly important for MOCO problems with more than two objectives.
Using these refined definitions, the following characterizations for MOCO problems hold: While the weakly supported non-dominated points lie on the boundary of the upper image, the supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces, and the unsupported non-dominated points lie in the interior of the upper image. However, these characterizations do not extend directly to general MOO problems. Hence, future research should aim to extend the characterization of supportedness beyond MOCO problems to general MOO contexts, addressing the gaps highlighted by this study. An approach towards a categorization of supportedness in the context of general multi-objective optimization has been discussed in Chlumsky-Harttmann, (2025).
Unsupported solutions may be reasonable compromise solutions and should thus not be neglected completely. Note that the difficulty in computing unsupported solutions arises in many integer and combinatorial optimization problems and is one reason for their computational complexity, in general (Ehrgott,, 2000; Figueira et al.,, 2017). One way to overcome this computational burden—at least to a certain degree—could be to determine unsupported solutions only in regions of the Pareto front that are not well represented by the set of supported non-dominated points.
Acknowledgements.
The authors thankfully acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project number 441310140.